In an A.P 2,5,8....., upto 50 terms
,and 3,5,7,9,... upto 60 terms,
find how many terms are identical?

Grade:12

1 Answers

bharat bajaj IIT Delhi
10 years ago
For first AP,
An = 2 + (x-1)3 = 3x - 1 where 0<x<51
For second AP,
Bn = 3 + (y-1)2 = 2y + 1 where 0<y<61
Now, for An = Bn
3x-1 = 2y + 1
3x - 2y = 2
where x and y are integers.
When x is an even no. , it satisfies this condition. Hence,
x=2, y = 2
x=4, y = 5
x=6, y = 8
.
.
.
.
x =40 , y= 59
Hence, 20 terms are identical.
Thanks
Bharat Bajaj
askiitians faculty
IIT Delhi